The Advanced Photon Source
a U.S. Department of Energy Office of Science User Facility

APS-UES Meeting Minutes - 12/01/2015

Attendees: Bruce Glagola, Wendy VanWingeren, Clay White, Bruno Fieramosca, Shane Flood, and Patti Pedergnana

Bruce Glagola

  • 24-ID has studies today from 1330-1730.  27-ID and 35-ID has parasitic studies.
  • CIS is having an annual maintenance this weekend.  This would affect Oracle databases, including ESAF and Training.  John MacLean is going to encourage all beamlines to approve their ESAFs early.
  • Today is Tech Tuesday; Bruno will attend 431, Clay 433, and Shane 438.
  • Patti will attend the Shutdown Planning meeting tomorrow morning.

Bruno Fieramosca

  • Bruno is having PC issues and is going to enter a HelpDesk.

Wendy VanWingeren

  • There are radioactive samples starting at 16-BMD and 10-IDB tomorrow.
  • 35-BM is getting a new Front End and Wendy discussed being RSSE for this work.

Clay White

  • 9-BM’s BLEPS is being disassembled now.  The disassembly of the beamline for installation of the new monochromator and movement of RSS components still has to be approved by the BSDRSC.  It will be discussed on Thursday.  The CCWP for disassembly of the beamline is in ICMS.  Disassembly can begin without BSDRSC approval.
  • There is an air leak at column 137 in the 436 area on the SR mezzanine.

Patti Pedergnana

  • BioCARS is having a funding review tomorrow.
  • HP-CAT is having their mono shutter cylinders rebuilt today, in preparation for the mono line upgrade.
  • 13-ID is having their Channel 1 Yokagawa transmitters bled today to help address the recurring PSS water flow below normal trips.